josiahcausey:

Which of the following inequalities matches the graph? (1 point) graph of an inequality with a dashed line through the points (0, 3) and (1, 9) with shading above the line −6x + y < 3 6x + y < 3 6x − y < −3 The correct inequality is not listed

Mercury:

start with the points (0,3) and (1,9) and write the equation first slope formula: m = (y2-y1) / (x2-x1) once you have your slope, you can plug slope and one of the points into the formula y = mx + b to find b. (I recommend using (0,3) as this makes the arithmetic easier) now you have your equation. "dashed" equation above the line implies all the values greater than the equation, so change y = mx + b to y > mx + b and find the corresponding inequality within the answer choices (if it is there). you may need re-arrange the equation to match one of the answer choices.